Grade 10General Physics

an electric iron has a rating of 250W,220V .calculate the current flowing through it and its resistance when in use

Nirupam Agarwal
9 Years agoGrade 10

1 Answer

Vikas TU
9 Years ago
 
We have given 
 
Power (P) = 250 W 
 
Potential Difference (V) = 220V 
 
As we realize that 
 
P = VI 
 
250 = 220 xI 
 
I = 25/22 
 
I = 1.36 A 
 
From the Ohms law 
 
V = IR 
 
220 = 1.36*R 
 
R = 220/1.36 
 
R = 161.76 ohm